OpenStudy (anonymous):

Write an equation and solve. Round to the nearest hundredth where necessary.19 is what percent of 40 ?

OpenStudy (anonymous):

(x/100)*40=19 do what did you get ?!

OpenStudy (anonymous):

so**

OpenStudy (anonymous):

i not sure 40=19 n,2.1% or 40=19n;40.4%

OpenStudy (anonymous):

:'( wrong equation ! 40*n=19

OpenStudy (anonymous):

so n=19/40

OpenStudy (anonymous):

so n=0,475 ! so th answer will be 0.475*100=47.5

OpenStudy (anonymous):

your are a grate !! thank you

OpenStudy (anonymous):

sorry ! 47.5 percent ------Yw anytime :) !

OpenStudy (anonymous):

your are got the right answer